OpenStudy (anonymous):

Which set of side lengths defines a right triangle? A. 4, 8, 9 B. 5, 9, 10 C. 6, 10, 13 D. 7, 24, 25

Nnesha (nnesha):

to test if given number defines a right triangle you have to use Pythagorean theorem \[\huge\rm a^2 +b^2 = c^2\] plug in a b and c value remember c is the longest side if right triangle if both sides are equal then yes those 3 lengths makes a right triangle

Nnesha (nnesha):

for example A) 4 . 8 , 9 c is 9 because it's the bigger number \[4^2 + 8^2 = 9^2\] solve it and let me know if both sides are equal or no :-)

OpenStudy (anonymous):

it equals 161 @Data_LG2

OpenStudy (anonymous):

wait i think i did it wrong im not sure

OpenStudy (anonymous):

ok its not equal i did it again

OpenStudy (anonymous):

and i tagged the wrong person xD

Nnesha (nnesha):

what did you get at left side and what u got at right side ?

Nnesha (nnesha):

yep try again

OpenStudy (anonymous):

for the left i got 80 and for the right 81

Nnesha (nnesha):

so both sides are not equal that's mean nope 4 , 8 and 9 are not makes right triangle now try answer choices

OpenStudy (anonymous):

i got D

Nnesha (nnesha):

both sides are equal ?

OpenStudy (anonymous):

yep

Nnesha (nnesha):

gO_Od to go! :-)

OpenStudy (anonymous):

thanks for the help :)

Nnesha (nnesha):

my pleasure :-)